A bus leaves from city A with a speed of 45km/h.After one hour a car leaves from the same point.With what speed does the car have to travell to reach te bus in less than three hours?
1
Expert's answer
2019-01-09T10:15:11-0500

Total time:

T=3+1=4 hours.

The distance travelled of the bus:

D=VT=(45)(4)=180 km.

The speed of the car:

v=D/t=180/3=60 km/h.